The perimeter of a rectangle is 130 in. Twice the length is five more than three times the width. Find the length and the width.
Answer by nyc_function(2741)

P = 2L + 2W
2L = 3w + 5
130 = 3w + 5 + 2w
130 = 5w + 5
130 - 5 = 5w
125 = 5w
125/5 = w
25 = w
We now replace w in the equation 2L = 3w + 5 with 25 and solve for L.
2L = 3(25) + 5
2L = 75 + 5
2L = 80
L = 80/2
L = 40
The width is 25 and the length is 40.
====================================
The Proof:
130 = 2(40) + 2(25)
130 = 80 + 50
130 = 130...It checks!!!
====================================
I received your reply. Sorry for the typo. The length is always longer than the width of a rectangle. So, the length = 40 and width = 25.